OpenStudy (anonymous):

how is [7,5,3] coplanar to [1,1,1] and [1,2,3]? thanks

OpenStudy (anonymous):

they are vectors btw

OpenStudy (amistre64):

define the plane .....

OpenStudy (amistre64):

cross 2 vectors to get the normal, then anchor it to one of the points .... tada, an equation to play with

OpenStudy (tkhunny):

If your matrices are up to speed, you can just GJ-Reduce and find the linear combination that you need for proof.

OpenStudy (anonymous):

does it start like this: 1=s1 1=s2 1=s3 ?

OpenStudy (amistre64):

not sure what s1s2s3 refer to

OpenStudy (amistre64):

have you gone over plane equations? or at best a dot product?

OpenStudy (anonymous):

[1,1,1]=s[1,2,3]

OpenStudy (amistre64):

thats not planar, thats parallel

OpenStudy (amistre64):

if they are coplanar, then there is one vector that will be perpendicular to all of them, let that vector be (a,b,c) and apply the dot product 7a+5b+3c=0 1a,+1b+1c=0 1a+2b+3c=0 a system of 3 equations in 3 unknowns like in any basic algebra class.

OpenStudy (anonymous):

here is the original question ... i actually got it wrong in a quiz but i only want to know how to solve it thats all

OpenStudy (amistre64):

well, there is still going to be some trial and error, but in this case i would develop the plane equation to test out the options, and not the matrix solutions

OpenStudy (amistre64):

tell me what you know of creating an equation of a plane from 2 vectors

OpenStudy (anonymous):

hmm.. getting confused

OpenStudy (amistre64):

the matrix process would create an infinite number of vectors as solution .... so thats not the route i would follow. what do you know about creating a plane equation?

OpenStudy (amistre64):

spose we define all the vectors that we can make from one of our given point, say from (1,1,1): the vector from (1,1,1) to some point (x,y,z) is just subtraction: (x-1,y-1,z-1) but this is all the vectors in a 3d space, we want to narrow them down to a plane .... using some vector (a,b,c) dotted to the vectors found, give us all the vectors that are perp to (a,b,c) a(x-1)+b(y-1)+c(z-1) = 0 finding that vector (a,b,c) is the key to developing the equation we need.

OpenStudy (amistre64):

it just so happens that finding the cross product of 2 vectors, defines a vector that is perp to both of them ... so we need to take the 2 given vectors, and cross them to find (a,b,c)